# arXiv:2603.00376v1[cs.AI] 27 Feb 2026

### 27 Feb 2026 v1

@@ II-A Cortical Columns and Reference Frames
Our NeuroAI approach is influenced by the work of Jeff Hawkins and Numenta [10], which was based on the discovery by Vernon Mountcastle of the columnar organization of the neocortex [19] and the theoretical neuroscience work on reference frames [1] [23]. Hawkins’ Thousand Brain Theory [11] suggests that Cortical Columns (CC) are the fundamental neocortical units that embody intelligence. CCs function as Reference Frames (RFs) that provide a sensorimotor modeling system. RFs record information through movement (sensorimotor learning), as the associated sensing agent moves through an environment, which could be physical or conceptual.
In Hawkins’ model, multiple RFs targeting different sensory modalities can interact with each other and seek consensus on the output through voting within and between sensory modalities. In contrast to current DNNs that separate training and inference, CCs that store and process information in RFs can support real-time inferencing, as well as online, concurrent, and continuous learning, and dynamically adapt to changes in sensory inputs from the environment. The continuous feedback loop of predict-sense-update within a reference frame effectively introduces an additional dimension of “memory” that remembers past observations and patterns. This “sequential” (or recurrent) behavior is missing in all feed-forward neural networks, e.g. CNNs and DNNs.
@@ IV-B NeuroHex Ring Encoding, Positional, Angular Arithmetics
Ring Encoding assigns a $RingIndex(RI)$ to every ring to identify itself relative to an arbitrary origin, and a $RingSpot$, which indicates the position of the coordinate on the designated ring. We define that the origin always starts at RIRI0”. A full ring can be subdivided into 6 segments with inclusive endpoints that correspond to the 6 wedges of a hexagon. Within the ring segment of a wedge, any $RingSpot$ will function as a relative $WedgeRingSpot$, effectively limiting its range from 0 - $RI$. An illustration of the basic components of NeuroHex is shown in Figure 4.
Due to the nature of hexagonal coordinates, the distance between two coordinates can be conveniently computed by taking the maximum of the absolute values of the differences of each cube coordinate. Any coordinate on a ring is also $RI$ away from the origin. See Figure 5 (left).
Directional computation in NeuroHex can be achieved via angular quantization. Connecting a line between every $WedgeRingSpot$ on the ring segment within a wedge to the origin, the lines effectively divide the 60° wedge into $RI$ angles, yielding quantization levels of $RI$. See Figure 5 (right). The entire 360° circle can be partitioned into $(6\times RI)$ equally spaced angles or quantized clockwise angular directions.
@@ IV-C3 Complex Shapes
By integrating simple Boolean logic into the shape-generation process, we can construct various geometries through additive overlap(AO)(AO) and subtractive overlap(SO)(SO). These two operations serve as complements: $SO$ isolates coordinates contained by one shape but excluded from the other, whereas $AO$ selects coordinates jointly bounded by both shapes. Some of the shapes generated include donuts (subtractive of two circles) and tear drops (additive of two circles). Using more than two foundational shapes further expands the geometric shapes that can be represented.
@@ IV-D NeuroHex Translation, Rotation, and Scaling of Shapes
Translation: In addition to allowing for easy origin translation, NeuroHex also enables efficient translation of magnitudemagnitude and quantized angle input to cubic coordinates. $RingIndex$ is equal to $(magnitude+1)$. The quantized angle allows for straightforward computation of direction: let the maximum per-wedge quantization level be a power of 2, encoded using $n$ bits in binary, the higher order $(n+2)$ to $n$ bits of the angle encoding can be used to index the 6 wedges, while the lower $n$ bits represent the wedge-local angular quantization. Determining the exact $WedgeRingSpot$ requires a conversion from relative quantization to integer length, which involves a single instance of logical shift and multiplication. Once $RingIndex$, WedgeIndexWedgeIndex, and $WedgeRingSpot$ have been derived, the target cubic coordinate resolution becomes trivial, since only the multiplexing selection of directions from the six wedges and potential sign reversal is needed without further arithmetic. This is illustrated in Figure 7.

### 03 Mar 2026 v2

# arXiv:2603.00376v2[cs.AI] 03 Mar 2026

@@ II-A Cortical Columns and Reference Frames
Our NeuroAI approach is influenced by the work of Jeff Hawkins and Numenta [10], which was based on the discovery by Vernon Mountcastle of the columnar organization of the neocortex [19] and the theoretical neuroscience work on reference frames [1] [23]. Hawkins’ Thousand Brain Theory [11] suggests that Cortical Columns ($C$) are the fundamental neocortical units that embody intelligence. $C$s function as Reference Frames (RFs) that provide a sensorimotor modeling system. RFs record information through movement (sensorimotor learning), as the associated sensing agent moves through an environment, which could be physical or conceptual.
In Hawkins’ model, multiple RFs targeting different sensory modalities can interact with each other and seek consensus on the output through voting within and between sensory modalities. In contrast to current DNNs that separate training and inference, $C$s that store and process information in RFs can support real-time inferencing, as well as online, concurrent, and continuous learning, and dynamically adapt to changes in sensory inputs from the environment. The continuous feedback loop of predict-sense-update within a reference frame effectively introduces an additional dimension of “memory” that remembers past observations and patterns. This “sequential” (or recurrent) behavior is missing in all feed-forward neural networks, e.g. CNNs and DNNs.
@@ IV-B NeuroHex Ring Encoding, Positional, Angular Arithmetics
Ring Encoding assigns a $RingIndex(RI)$ to every ring to identify itself relative to an arbitrary origin, and a RingSpotRingSpot, which indicates the position of the coordinate on the designated ring. We define that the origin always starts at RIRI0”. A full ring can be subdivided into 6 segments with inclusive endpoints that correspond to the 6 wedges of a hexagon. Within the ring segment of a wedge, any RingSpotRingSpot will function as a relative WedgeRingSpotWedgeRingSpot, effectively limiting its range from 0 - RIRI. An illustration of the basic components of NeuroHex is shown in Figure 4.
Due to the nature of hexagonal coordinates, the distance between two coordinates can be conveniently computed by taking the maximum of the absolute values of the differences of each cube coordinate. Any coordinate on a ring is also RIRI away from the origin. See Figure 5 (left).
Directional computation in NeuroHex can be achieved via angular quantization. Connecting a line between every WedgeRingSpotWedgeRingSpot on the ring segment within a wedge to the origin, the lines effectively divide the 60° wedge into RIRI angles, yielding quantization levels of RIRI. See Figure 5 (right). The entire 360° circle can be partitioned into $(6\times RI)$ equally spaced angles or quantized clockwise angular directions.
@@ IV-C3 Complex Shapes
By integrating simple Boolean logic into the shape-generation process, we can construct various geometries through additive overlap(AO)(AO) and subtractive overlap(SO)(SO). These two operations serve as complements: SOSO isolates coordinates contained by one shape but excluded from the other, whereas AOAO selects coordinates jointly bounded by both shapes. Some of the shapes generated include donuts (subtractive of two circles) and tear drops (additive of two circles). Using more than two foundational shapes further expands the geometric shapes that can be represented.
@@ IV-D NeuroHex Translation, Rotation, and Scaling of Shapes
Translation: In addition to allowing for easy origin translation, NeuroHex also enables efficient translation of magnitudemagnitude and quantized angle input to cubic coordinates. RingIndexRingIndex is equal to $(magnitude+1)$. The quantized angle allows for straightforward computation of direction: let the maximum per-wedge quantization level be a power of 2, encoded using $n$ bits in binary, the higher order $(n+2)$ to $n$ bits of the angle encoding can be used to index the 6 wedges, while the lower $n$ bits represent the wedge-local angular quantization. Determining the exact WedgeRingSpotWedgeRingSpot requires a conversion from relative quantization to integer length, which involves a single instance of logical shift and multiplication. Once RingIndexRingIndex, WedgeIndexWedgeIndex, and WedgeRingSpotWedgeRingSpot have been derived, the target cubic coordinate resolution becomes trivial, since only the multiplexing selection of directions from the six wedges and potential sign reversal is needed without further arithmetic. This is illustrated in Figure 7.